Toby company sells mobile phones at $ 80 and profits were $ 9,600 a week.
The company is considering reducing prices by 20% to increase its market share. Estimated price elasticity of demand is 1.6.
Required
Estimate the effect of price cuts on company revenues.
If a competing company reduces the price by 20% compared to Toby , what effect does this have on Tobi, assuming that the elasticity of the cross demand is 0.8.
1
Expert's answer
2020-01-03T09:32:42-0500

If price elasticity of demand is 1.6, it means that quantity demanded rises by 20×1.6=32%. Then, company revenues change is 1.32Q1×0.8P1-Q1×P1=0.05×Q1P1 (increases by 5%) = 0.05×9,600=0,480 $

If a competing company reduces the price by 20% and cross elasticity is 0.8, it means that quantity demanded decreases by 20×0.8=16%. Then, company revenues change is 0.84Q1×P1-Q1×P1=-0.16Q1P1 (decreases by 16%)=- 0.16×9,600=-1,530 $